# A Unified View of Optimal Kernel Hypothesis Testing

### v1

@@ 6.3 Uniform power against alternatives separated in L2 metric
The KSD test with some specific kernel bandwidth depending on the unknown Sobolev smoothness is guaranteed to be powerful when (Section A.7)
$\big{\|}\big{(}\nabla\log p-\nabla\log q\big{)}\,q\,\big{\|}_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{N}}\right)^{2s/(4s+5d)}$ (49)
@@
To avoid the dependence on the unknown Sobolev smoothness $s$ for practical uses, one can rely on aggregating over a collection of kernel bandwidths independent of $s$ (multiple testing), to obtain the uniform separation rate (Section A.9)
$\big{\|}\big{(}\nabla\log p-\nabla\log q\big{)}\,q\,\big{\|}_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{N/\log(\log(N))}}\right)^{2s/(4s+5d)}.$ (50)
@@
Using a KSD estimator with design ${\mathcal{D}}$, the goodness-of-fit test with kernel bandwidth depending on $s$, the unknown Sobolev smoothness, controls the type II error by $\beta$ for alternatives satisfying (Section A.8)
$\big{\|}\big{(}\nabla\log p-\nabla\log q\big{)}\,q\,\big{\|}_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{|{\mathcal{D}}|/N}}\right)^{2s/(4s+5d)}.$ (51)
Again, if a complete statistic is used (i.e., $|{\mathcal{D}}|\asymp N^{2}$), then this is the same rate as Equation 49 in the standard framework. There is a trade-off between efficiency and power in terms of speed of the uniform separation rate: when the complexity $|{\mathcal{D}}|$ decreases, the set of detectable alternatives shrinks, until it is finally empty for linear tests with $|{\mathcal{D}}|\asymp N$.
When relying on aggregation over various kernel bandwidths independently of the unknown Sobolev smoothness (multiple testing), the resulting test is powerful provided that (Section A.9)
$\big{\|}\big{(}\nabla\log p-\nabla\log q\big{)}\,q\,\big{\|}_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{\big{(}|{\mathcal{D}}|/N\big{)}\big{/}\log\!\big{(}\!\log(|{\mathcal{D}}|/N)\big{)}}}\right)^{2s/(4s+5d)}$ (52)
@@ 7 Open problems for future work
    $L^{2}$ separation upper bound for the HSIC test which also holds for smoothness $s\in\big{(}0,(d_{\mathcal{X}}+d_{\mathcal{Y}})/4\big{)}$ (e.g., Equation 36).
@@
    $L^{2}$ separation lower bounds under differential privacy constraint (e.g., Equations 25, 26, 27, 40, 41 and 42).
@@
    $L^{2}$ and kernel separation lower bounds for efficient testing under computational complexity constraint (e.g., Equations 16, 31, 47, 23, 38 and 51).
@@
    $L^{2}$ separation lower bounds, and improved upper bounds, for goodness-of-fit testing (e.g., Equation 49).
@@
    Kernel separation rates for normalised pooling (e.g., Equations 15, 30 and 46 for unnormalised, Biggs et al., 2023, Theorem 3 for normalised).
@@
    Kernel pooling and aggregation procedure for differentially private tests with noise scaled independently of the number of kernels (see discussion below Equation 18).
@@
    KSD private and robust test constructions, and uniform separation power guarantees (see discussions in Section 6.1).
@@ Appendix A Proof sketches
  • Section A.1: Proof sketch of kernel separation (V/U-statistics).
  • Section A.2: Proof sketch of differentially private kernel separation (V-statistics).
  • Section A.3: Proof sketch of robust kernel separation (V-statistics).
  • Section A.4: Proof sketch of pooled kernel separation (V/U-statistics).
  • Section A.5: Proof sketch of efficient kernel separation (B-statistics).
  • Section A.6: Proof sketch of pooled efficient kernel separation (B-statistics).
  • Section A.7: Proof sketch of $L^{2}$ separation (U-statistics).
  • Section A.8: Proof sketch of efficient $L^{2}$ separation (incomplete U-statistics).
  • Section A.9: Proof sketch of aggregated $L^{2}$ separation (U-statistics).
  • Section A.10: Proof sketch of aggregated efficient $L^{2}$ separation (incomplete U-statistics).
  • Section A.11: Proof sketch of differentially private $L^{2}$ separation (U-statistics).
In general, the kernel separation results (Sections A.1, A.2, A.5, A.3, A.4 and A.6) hold naturally when using the square-rooted V-statistic $\sqrt{V_{k}}$ with kernel $k$, and the $L^{2}$ separation results (Sections A.7, A.8, A.9, A.10 and A.11) hold naturally when using the unbiased U-statistic $U_{k}$ with kernel $k$. The classical kernel separation result (Section A.1) can also be proved for U-statistics by leveraging their relation to V-statistics, and the efficient kernel separation result (Section A.5) holds only for block B-statistics (not other incomplete U-statistics) because these can be related to a non-negative block incomplete V-statistic version. Meanwhile, the efficient (aggregated) $L^{2}$ separation results (Sections A.8 and A.10) hold for any incomplete U-statistics.
@@ A.1 Proof sketch of kernel separation
We detail the proof structure of the kernel separation results: Equation 13 proved in Kim and Schrab (2023, Theorem 7), Equation 28 proved in Kim and Schrab (2023, Theorem 12), and Equation 44 proved here.
@@
Consider the statistic $\sqrt{V_{0}}$, where $V_{0}\coloneqq V_{k}$ is a V-statistic (e.g., Kim and Schrab, 2023, Equations 7, 19 and 52) for some fixed kernel $k$, $\sqrt{V_{0}}$ is an estimate of some kernel discrepancy ${\mathrm{Kdisc}}$ (e.g., MMD, HSIC, KSD) for which ${\mathrm{Kdisc}}=0$ characterises the null. We also consider bootstrapped statistics $\sqrt{V_{1}},\dots,\sqrt{V_{B}}$ (either via permutations or wild bootstrap, Section 1.2), and the $(1-\alpha)$-quantile $q_{1-\alpha}$ of $\sqrt{V_{0}},\dots,\sqrt{V_{B}}$. The aim is to bound the probability of type II error $\mathbb{P}(T_{0}\leq q_{1-\alpha})$ by $\beta$ provided that the discrepancy ${\mathrm{Kdisc}}$ is greater than some rate to be determined. For this, we need two results: some exponential concentration bounds for the original statistic and for the bootstrapped statistic.
@@
$\mathbb{P}\Big{(}\big{|}\sqrt{V_{0}}-{\mathrm{Kdisc}}\big{|}>\widetilde{C}\,N^{-1/2}+t\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \exp(-C\,t^{2}N)$ (53)
@@
$\big{|}\sqrt{V_{0}}-{\mathrm{Kdisc}}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{1}\sqrt{\frac{1}{N}\log\biggl{(}\frac{1}{\beta}\biggr{)}}$ (54)
with probability at least $21-\beta/$ 2. Such statistic concentration results can be found in Kim and Schrab (2023, Lemma 13) for MMD, and in Kim and Schrab (2023, Lemma 14) for HSIC. The KSD V-statistic case can be derived from the KSD U-statistic case (holding by Hoeffding’s inequality, Hoeffding, 1963, Equation 5.7) using the two facts that $U_{\mathrm{KSD}}(N-1)/N\leq V_{\mathrm{KSD}}\leq K/N+U_{\mathrm{KSD}}(N-1)/N$, where $K$ is a bound on the kernel, and that $\big{|}V_{0}-{\mathrm{Kdisc}}^{2}\big{|}=\big{|}\sqrt{V_{0}}-{\mathrm{Kdisc}}\big{|}\big{|}\sqrt{V_{0}}+{\mathrm{Kdisc}}\big{|}\leq 2K\big{|}\sqrt{V_{0}}-{\mathrm{Kdisc}}\big{|}$. The U-statistics and V-statistics for MMD and HSIC can also be expressed in terms of each other (Kim and Schrab, 2023, Appendix E.13 and Lemma 22).
@@
$\mathbb{P}\Big{(}\sqrt{T_{b}}>\widetilde{C}\,N^{-1/2}+t\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \exp(-C\,t^{2}N)$ (55)
for all $t>0$, which leads to an upper bound on the quantile [^9] The concentration inequality results in a bound on the quantile $q^{\infty}_{1-\alpha}$ obtained with infinitely many bootstrapped statistics. Kim and Schrab (2023, Lemma 21) then guarantees that this translates to a bound on the quantile $q_{1-\alpha}$ obtained with a finite number of bootstrapped statistics. For example, as illustrated in Kim and Schrab (2023, Equation 63), if this number is larger than $6\log(2/\beta)/\alpha$ then $q_{1-\alpha}\leq q^{\infty}_{1-\alpha/6}$ and the bound directly applies to the quantile with finitely many bootstrapped statistics noting that $\log(6/\alpha)\lesssim\log(1/\alpha)$ with $\log(1/\alpha)\geq 1$. So, even though the quantile bound (which holds with probability at least $21-\beta/$ 2) does not at first appear to depend on $\beta$, this dependence is hidden in the condition on the number of bootstrapped statistics. This reasoning holds throughout the proofs of this section and we do not explicitly mention the condition on the number of bootstrapped statistics (e.g., greater than $6\log(2/\beta)/\alpha$) every time.
$q_{1-\alpha}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{2}\sqrt{\frac{1}{N}\log\biggl{(}\frac{1}{\alpha}\biggr{)}},$ (56)
holding with probability at least $21-\beta/$ 2. This concentration bound can be obtained when using permutations for MMD and HSIC (Kim and Schrab, 2023, Lemmas 10 and 12), and when using wild bootstrap for MMD, HSIC and KSD (Rademacher chaos concentration of de la Peña and Giné, 1999b, Corollary 3.2.6).
@@
${\mathrm{Kdisc}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}},$ (58)
@@
Consider the U-statistic $U_{0}\coloneqq U_{k}$ for some fixed kernel $k$ (e.g., Schrab, 2025b, Equations 12, 26 and 52), which is an estimate of some kernel squared discrepancy ${\mathrm{Kdisc}}^{2}$ (e.g., MMD2, HSIC2, KSD2) for which ${\mathrm{Kdisc}}=0$ characterises the null. We also consider bootstrapped statistics ${U_{1}},\dots,{U_{B}}$ (either via permutations or wild bootstrap, Section 1.2), and the $(1-\alpha)$-quantile $q_{1-\alpha}$ of ${U_{0}},\dots,{U_{B}}$. Again, the aim is to bound the probability of type II error $\mathbb{P}(T_{0}\leq q_{1-\alpha})$ by $\beta$ provided that the discrepancy ${\mathrm{Kdisc}}$ is greater than some rate to be determined. This can be derived by first using inequalities linking the U-statistic and V-statistic, and then using using some exponential concentration bounds for the bootstrapped U-statistic.
@@
$\mathbb{P}\Big{(}\big{|}{U_{0}}-{\mathrm{Kdisc}}^{2}\big{|}>t\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \exp(-C\,t^{2}N)$ (59)
@@
$\big{|}{U_{0}}-{\mathrm{Kdisc}}^{2}\big{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{1}\sqrt{\frac{1}{N}\log\biggl{(}\frac{1}{\beta}\biggr{)}}$ (60)
with probability at least $21-\beta/$ 2. For MMD, HSIC and KSD, this holds by Hoeffding’s inequality (Hoeffding, 1963, Equation 5.7) provided that the kernels are bounded.
@@
$U_{0}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ V_{0}-\frac{2K}{N}$ (61)
@@
$\mathbb{P}\Big{(}{U_{b}}>t\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \exp(-C\,tN)$ (62)
@@
$q_{1-\alpha}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{3}{\frac{1}{N}\log\biggl{(}\frac{1}{\alpha}\biggr{)}},$ (63)
holding with probability at least $21-\beta/$ 2. For permutations (MMD and HSIC), this concentration bound for permuted U-statistics holds by (Kim et al., 2022, Theorems 6.1, 6.2 & 6.3). For the wild bootstrap method (MMD, HSIC and KSD), the Rademacher chaos concentration of de la Peña and Giné (1999b, Corollary 3.2.6) gives the desired quantile bound. See also the details of deriving the quantile bounds in Section A.7, Schrab et al. (2023, Proposition 4), Schrab et al. (2022a, Theorem 3.1), and Schrab et al. (2022b, Lemma 2).
@@
${\mathrm{Kdisc}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}},$ (65)

### v2

@@ 6.3 Uniform power against alternatives separated in L2 metric
The KSD test with some specific kernel bandwidth depending on the unknown Sobolev smoothness is guaranteed to be powerful when (Section˜A.7)
$\big\|\big(\nabla\log p-\nabla\log q\big)\,q\,\big\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{N}}\right)^{2s/(4s+5d)}$ (49)
@@
To avoid the dependence on the unknown Sobolev smoothness $s$ for practical uses, one can rely on aggregating over a collection of kernel bandwidths independent of $s$ (multiple testing), to obtain the uniform separation rate (Section˜A.9)
$\big\|\big(\nabla\log p-\nabla\log q\big)\,q\,\big\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{N/\log(\log(N))}}\right)^{2s/(4s+5d)}.$ (50)
@@
Using a KSD estimator with design $\mathcal{D}$}, the goodness-of-fit test with kernel bandwidth depending on $s$, the unknown Sobolev smoothness, controls the type II error by $\beta$ for alternatives satisfying (Section˜A.8)
$\big\|\big(\nabla\log p-\nabla\log q\big)\,q\,\big\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{|{\mathcal{D}}|/N}}\right)^{2s/(4s+5d)}.$ (51)
Again, if a complete statistic is used (i.e., $|{\mathcal{D}}|\asymp N^{2}$), then this is the same rate as Equation˜49 in the standard framework. There is a trade-off between efficiency and power in terms of speed of the uniform separation rate: when the complexity $|{\mathcal{D}}|$ decreases, the set of detectable alternatives shrinks, until it is finally empty for linear tests with $|{\mathcal{D}}|\asymp N$.
When relying on aggregation over various kernel bandwidths independently of the unknown Sobolev smoothness (multiple testing), the resulting test is powerful provided that (Section˜A.9)
$\big\|\big(\nabla\log p-\nabla\log q\big)\,q\,\big\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\sqrt{\big(|{\mathcal{D}}|/N\big)\big/\log\!\big(\!\log(|{\mathcal{D}}|/N)\big)}}\right)^{2s/(4s+5d)}$ (52)
@@ 7 Open problems for future work
    $L^{2}$ separation upper bound for the HSIC test which also holds for smoothness $s\in\big(0,(d_{\mathcal{X}}+d_{\mathcal{Y}})/4\big)$ (e.g., Equation˜36).
@@
    $L^{2}$ separation lower bounds under differential privacy constraint (e.g., Equations˜25, 26, 27, 40, 41 and 42).
@@
    $L^{2}$ and kernel separation lower bounds for efficient testing under computational complexity constraint (e.g., Equations˜16, 31, 47, 23, 38 and 51).
@@
    $L^{2}$ separation lower bounds, and improved upper bounds, for goodness-of-fit testing (e.g., Equation˜49).
@@
    Kernel separation rates for normalised pooling (e.g., Equations˜15, 30 and 46 for unnormalised, Biggs et al., 2023, Theorem 3 for normalised).
@@
    Kernel pooling and aggregation procedure for differentially private tests with noise scaled independently of the number of kernels (see discussion below Equation˜18).
@@
    KSD private and robust test constructions, and uniform separation power guarantees (see discussions in Section˜6.1).
@@ Appendix A Proof sketches
  • Section˜A.1: Proof sketch of kernel separation (V/U-statistics).
  • Section˜A.2: Proof sketch of differentially private kernel separation (V-statistics).
  • Section˜A.3: Proof sketch of robust kernel separation (V-statistics).
  • Section˜A.4: Proof sketch of pooled kernel separation (V/U-statistics).
  • Section˜A.5: Proof sketch of efficient kernel separation (B-statistics).
  • Section˜A.6: Proof sketch of pooled efficient kernel separation (B-statistics).
  • Section˜A.7: Proof sketch of $L^{2}$ separation (U-statistics).
  • Section˜A.8: Proof sketch of efficient $L^{2}$ separation (incomplete U-statistics).
  • Section˜A.9: Proof sketch of aggregated $L^{2}$ separation (U-statistics).
  • Section˜A.10: Proof sketch of aggregated efficient $L^{2}$ separation (incomplete U-statistics).
  • Section˜A.11: Proof sketch of differentially private $L^{2}$ separation (U-statistics).
In general, the kernel separation results (Sections˜A.1, A.2, A.5, A.3, A.4 and A.6) hold naturally when using the square-rooted V-statistic $\sqrt{V_{k}}$ with kernel $k$, and the $L^{2}$ separation results (Sections˜A.7, A.8, A.9, A.10 and A.11) hold naturally when using the unbiased U-statistic $U_{k}$ with kernel $k$. The classical kernel separation result (Section˜A.1) can also be proved for U-statistics by leveraging their relation to V-statistics, and the efficient kernel separation result (Section˜A.5) holds only for block B-statistics (not other incomplete U-statistics) because these can be related to a non-negative block incomplete V-statistic version. Meanwhile, the efficient (aggregated) $L^{2}$ separation results (Sections˜A.8 and A.10) hold for any incomplete U-statistics.
@@ A.1 Proof sketch of kernel separation
We detail the proof structure of the kernel separation results: Equation˜13 proved in Kim and Schrab (2023, Theorem 7), Equation˜28 proved in Kim and Schrab (2023, Theorem 12), and Equation˜44 proved here.
@@
Consider the statistic $\sqrt{V_{0}}$, where $V_{0}\coloneqq V_{k}$ is a V-statistic (e.g., Kim and Schrab, 2023, Equations 7, 19 and 52) for some fixed kernel $k$, $\sqrt{V_{0}}$ is an estimate of some kernel discrepancy $\mathrm{Kdisc}$} (e.g., MMD, HSIC, KSD) for which ${\mathrm{Kdisc}}=0$ characterises the null. We also consider bootstrapped statistics $\sqrt{V_{1}},\dots,\sqrt{V_{B}}$ (either via permutations or wild bootstrap, Section˜1.2), and the $(1-\alpha)$-quantile $q_{1-\alpha}$ of $\sqrt{V_{0}},\dots,\sqrt{V_{B}}$. The aim is to bound the probability of type II error $\mathbb{P}(T_{0}\leq q_{1-\alpha})$ by $\beta$ provided that the discrepancy $\mathrm{Kdisc}$} is greater than some rate to be determined. For this, we need two results: some exponential concentration bounds for the original statistic and for the bootstrapped statistic.
@@
$\mathbb{P}\Big(\big|\sqrt{V_{0}}-{\mathrm{Kdisc}}\big|>\widetilde{C}\,N^{-1/2}+t\Big)\penalty 10000\ \leq\penalty 10000\ \exp(-C\,t^{2}N)$ (53)
@@
$\big|\sqrt{V_{0}}-{\mathrm{Kdisc}}\big|\penalty 10000\ \leq\penalty 10000\ C_{1}\sqrt{\frac{1}{N}\log\biggl(\frac{1}{\beta}\biggr)}$ (54)
with probability at least $1-\beta/2$. Such statistic concentration results can be found in Kim and Schrab (2023, Lemma 13) for MMD, and in Kim and Schrab (2023, Lemma 14) for HSIC. The KSD V-statistic case can be derived from the KSD U-statistic case (holding by Hoeffding’s inequality, Hoeffding, 1963, Equation 5.7) using the two facts that $U_{\mathrm{KSD}}(N-1)/N\leq V_{\mathrm{KSD}}\leq K/N+U_{\mathrm{KSD}}(N-1)/N$, where $K$ is a bound on the kernel, and that $\big|V_{0}-{\mathrm{Kdisc}}^{2}\big|=\big|\sqrt{V_{0}}-{\mathrm{Kdisc}}\big|\big|\sqrt{V_{0}}+{\mathrm{Kdisc}}\big|\leq 2K\big|\sqrt{V_{0}}-{\mathrm{Kdisc}}\big|$. The U-statistics and V-statistics for MMD and HSIC can also be expressed in terms of each other (Kim and Schrab, 2023, Appendix E.13 and Lemma 22).
@@
$\mathbb{P}\Big(\sqrt{T_{b}}>\widetilde{C}\,N^{-1/2}+t\Big)\penalty 10000\ \leq\penalty 10000\ \exp(-C\,t^{2}N)$ (55)
for all $t>0$, which leads to an upper bound on the quantile [^9] The concentration inequality results in a bound on the quantile $q^{\infty}_{1-\alpha}$ obtained with infinitely many bootstrapped statistics. Kim and Schrab (2023, Lemma 21) then guarantees that this translates to a bound on the quantile $q_{1-\alpha}$ obtained with a finite number of bootstrapped statistics. For example, as illustrated in Kim and Schrab (2023, Equation 63), if this number is larger than $6\log(2/\beta)/\alpha$ then $q_{1-\alpha}\leq q^{\infty}_{1-\alpha/6}$ and the bound directly applies to the quantile with finitely many bootstrapped statistics noting that $\log(6/\alpha)\lesssim\log(1/\alpha)$ with $\log(1/\alpha)\geq 1$. So, even though the quantile bound (which holds with probability at least $1-\beta/2$) does not at first appear to depend on $\beta$, this dependence is hidden in the condition on the number of bootstrapped statistics. This reasoning holds throughout the proofs of this section and we do not explicitly mention the condition on the number of bootstrapped statistics (e.g., greater than $6\log(2/\beta)/\alpha$) every time.
$q_{1-\alpha}\penalty 10000\ \leq\penalty 10000\ C_{2}\sqrt{\frac{1}{N}\log\biggl(\frac{1}{\alpha}\biggr)},$ (56)
holding with probability at least $1-\beta/2$. This concentration bound can be obtained when using permutations for MMD and HSIC (Kim and Schrab, 2023, Lemmas 10 and 12), and when using wild bootstrap for MMD, HSIC and KSD (Rademacher chaos concentration of de la Peña and Giné, 1999b, Corollary 3.2.6).
@@
${\mathrm{Kdisc}}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}},$ (58)
@@
Consider the U-statistic $U_{0}\coloneqq U_{k}$ for some fixed kernel $k$ (e.g., Schrab, 2025b, Equations 12, 26 and 52), which is an estimate of some kernel squared discrepancy ${\mathrm{Kdisc}}^{2}$ (e.g., MMD2, HSIC2, KSD2) for which ${\mathrm{Kdisc}}=0$ characterises the null. We also consider bootstrapped statistics ${U_{1}},\dots,{U_{B}}$ (either via permutations or wild bootstrap, Section˜1.2), and the $(1-\alpha)$-quantile $q_{1-\alpha}$ of ${U_{0}},\dots,{U_{B}}$. Again, the aim is to bound the probability of type II error $\mathbb{P}(T_{0}\leq q_{1-\alpha})$ by $\beta$ provided that the discrepancy $\mathrm{Kdisc}$} is greater than some rate to be determined. This can be derived by first using inequalities linking the U-statistic and V-statistic, and then using using some exponential concentration bounds for the bootstrapped U-statistic.
@@
$\mathbb{P}\Big(\big|{U_{0}}-{\mathrm{Kdisc}}^{2}\big|>t\Big)\penalty 10000\ \leq\penalty 10000\ \exp(-C\,t^{2}N)$ (59)
@@
$\big|{U_{0}}-{\mathrm{Kdisc}}^{2}\big|\penalty 10000\ \leq\penalty 10000\ C_{1}\sqrt{\frac{1}{N}\log\biggl(\frac{1}{\beta}\biggr)}$ (60)
with probability at least $1-\beta/2$. For MMD, HSIC and KSD, this holds by Hoeffding’s inequality (Hoeffding, 1963, Equation 5.7) provided that the kernels are bounded.
@@
$U_{0}\penalty 10000\ \geq\penalty 10000\ V_{0}-\frac{2K}{N}$ (61)
@@
$\mathbb{P}\Big({U_{b}}>t\Big)\penalty 10000\ \leq\penalty 10000\ \exp(-C\,tN)$ (62)
@@
$q_{1-\alpha}\penalty 10000\ \leq\penalty 10000\ C_{3}{\frac{1}{N}\log\biggl(\frac{1}{\alpha}\biggr)},$ (63)
holding with probability at least $1-\beta/2$. For permutations (MMD and HSIC), this concentration bound for permuted U-statistics holds by (Kim et al., 2022, Theorems 6.1, 6.2 & 6.3). For the wild bootstrap method (MMD, HSIC and KSD), the Rademacher chaos concentration of de la Peña and Giné (1999b, Corollary 3.2.6) gives the desired quantile bound. See also the details of deriving the quantile bounds in Section˜A.7, Schrab et al. (2023, Proposition 4), Schrab et al. (2022a, Theorem 3.1), and Schrab et al. (2022b, Lemma 2).
@@
${\mathrm{Kdisc}}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}},$ (65)